A bag contains 6 Red, 4 Blue, and 10 Green marbles.
Two marbles are drawn, with replacement.
What's the probability of drawing Two Red marbles is a row?

Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

Sampling without replacement
first one is 6/20 probability red
second is 5/19 red
that product (they are multiplied) is 30/380 or 3/38=0.0789

Answer by greenestamps(13367)



The problem says sampling WITH replacement; the probability of red on each draw is 6/10 = 3/5.

P(red,red) = (3/5)(3/5) = 9/25

Convert to decimal or percent if required.